When the price of good X falls from . 10 to . 9, the demand for good Y
increase from 20 Kg. to 25 Kg.
a) What is the cross elasticity of demand of good Y for good X?
b) Are goods X and Y compliments or substitutes?
1
Expert's answer
2012-02-14T10:46:28-0500
Cross elasticity:
Ey,x= %change in quantity demanded of product Y/ %change in price of product X
%change in quantity demanded of product Y = (25/20 -1) *100% = 25%
%change in price of product X = (9/10-1)*100%= -10%
Ey,x= 25%/(-10%) = -2.5
If the cross elasticity of demand is negative than goods are complements. When price on X decreases the Demand for Y increases